Three charges, q1 = +1.90 10-9 C, q2 = -2.91 10-9 C, and q3 = +1.10 10-9 C, are located on the x-axis at x1 = 0, x2 = 10.0 cm, and x3 = 20.0 cm. Find the resultant force on q3.
Answer #1

Given 1.9 * 10-9C

y-_-2.91 * 10-9C

31.1 10-90

d1 = 0.2m

d2 = 20cm-10cm 10cm 0.1m

92-2.91nC 3 1.1nC 10cm 10cm F1

F_{1}=kq_{1}q_{3}/d_{1}^{2}

Fi = 9 * 109 * 1.9 * 10-9C * 1.1 * 10-9C/(0.2771)

F1 9 101.9 10100.22

F, 1.881 * 10-8/0.22

F1 4.702510N towards right

====================

F_{2}=kq_{2}q_{3}/d_{2}^{2}

F2-9 * 10% * 2.91 * 10-9C * 1.1 * 10-9C/(0.1m)2

F2-9 * 109 * 2.91 * 10-9 * 1.1 * 10-9/0.12

F_{2}=2.8809*10^{-8}/0.1^{2}

F_{2}=2.8809*10^{-6}N towards left

========================

Resultant Force ,F=F_{1}-F_{2}

F=4.7025*10^{-7}N-2.8809*10^{-6}N

ANSWER: F=-2.41065*10^{-6}N (negative sign shows direction as towards left)

======================================================================